OpenStudy (anonymous):

The Process: a. List of Sporting Balls: 1. Soccer Ball 2. Baseball 3. Basket Ball 4. Bowling Ball 5. Tennis Ball 6. Volleyball 7. Water Polo Ball. b. Ball Radius Diameter Circumference Surface Area Volume Ex: Ball x. Soccer Ball: R= d/2 = 8.75 in. Diameter of a Soccer Ball =2r =2 times 8.75 =17.5 in. D=Diameter, D=17.5 inches. A Soccer Ball is 27-28 inches in circumference. Circumference is pi times diameter, or 2 times pi times radius. Radius = circumference divided by (2 times pi). C= Circumference, C= 27-28 inches. Surface area of a

OpenStudy (anonymous):

The Process: a. List of Sporting Balls: 1. Soccer Ball 2. Baseball 3. Basket Ball 4. Bowling Ball 5. Tennis Ball 6. Volleyball 7. Water Polo Ball. b. Ball Radius Diameter Circumference Surface Area Volume Ex: Ball x. Soccer Ball: R= d/2 = 8.75 in. Diameter of a Soccer Ball =2r =2 times 8.75 =17.5 in. D=Diameter, D=17.5 inches. A Soccer Ball is 27-28 inches in circumference. Circumference is pi times diameter, or 2 times pi times radius. Radius = circumference divided by (2 times pi). C= Circumference, C= 27-28 inches. Surface area of a Soccer Ball =4 pi r squared =4 times 3.14 times 8.75 squared =232.2 square inches. Surface Area =SA, SA= 232.2 square inches Volume of a Soccer Ball =(4/3)(pi)(radius)3 cubic units, 2 times pi times r = 0.6858 r=0.109m, V=(4/3) times pi times (0.109)^3 =V V=0.005425m^3 =0.21 cubic inches. Baseball: Radius of a Baseball = d/2 R =1.77 inches. Diameter of a Baseball =2r D= 3.54 inches. Circumference of a Baseball = pi times diameter =C =Circumference, C=9-9.25. Surface Area of a Baseball SA =4 times pi times radius squared SA=25.8-27.2 square inches. Volume of a Baseball V =4 times pi times r cubed over 3 V=12.31 cubic inches. Basket Ball: Radius of a Basketball = d/2 R =4.7 inches. Diameter of Basketball =2r D=9.39 inches Circumference of a Basketball =pi times diameter C=29.5-30 inches. Surface Area of a Basketball = 4 times pi times radius squared SA= 286.5 square inches Volume of Basketball =4 times pi times r cubed over 3 V=433.526 cubic inches. Bowling Ball: Radius of a Bowling Ball =d/2 =a tenpin bowling ball, the radius is 4.25 inches. R= Radius, R= 4.25 inches. Diameter of a Bowling Ball =2r =2 times 4.25 = 8.5 D=Diameter, D= 8.5 inches. Circumference of a Bowling Ball =pi times diameter = 3.14 times 8.5= = about 53.4. C= Circumference, C= 53.4 inches Surface Area of a Bowling Ball = 4 times pi times radius squared = 4 times 3.14 times 4.25 squared, SA= 226.9 squared inches Volume of a Bowling Ball =4 times pi times r cubed over 3 =4 * 3.1416 *4.5*4.5*4.5 /3 = 381.7 V= Volume V= 381.7 cubic inches. Tennis Ball: Radius of a Tennis Ball =d/2 R = 2.57-2.7. Diameter of a Tennis Ball =2r D=Diameter,,D = 5.4. Circumference of a Tennis Ball = pi times diameter C=Circumference, C = 3.14 times 5.4 = about 16.9 C= 16.9 inches. Surface Area of a Tennis Ball = 4 times pi times radius squared =4 times 3.14 times 2.7 =33.9 SA= Surface Area, SA= 33.9 squared inches. Volume of a Tennis Ball = 4 times pi times r cubed over 3 = 4 times 3.14 times 2.7 cubed = 247.2 =V V= Volume, V= 247.2 cubic inches. Volley Ball: Radius of a Volley Ball = d/2 =25.6-26.4. R= Radius, R = 25.6-26. Diameter of a Volley Ball = 2r D =Diameter, D = approx. 52 inches. Circumference of a Volley Ball = pi times diameter = 3.14 times 26 = 81.64 C=Circumference, C= 81.64 inches. Surface Area of a Volley Ball = 4 times pi times radius squared = 4 times 3.14 times 26 = approx. 326.5. SA = Surface Area, SA = 326.5 squared inches. Volume of a Volley Ball = 4 times pi times r cubed over 3 =4 times 3.14 times 26 times 26 times 26 over 3 =73 584.85. V=Volume, V=73,584.85 cubic inches. Water Polo Ball Radius of a Water Polo Ball = d/2 = 8.6. R=Radius R=8.6 inches. Diameter of a Water Polo Ball =2r D=Diameter D = 17.2 inches. Circumference of a Water Polo Ball = pi times diameter = 3.14 times 8.6 = 27 C=Circumference C=27 inches. Surface Area of a Volley Ball = 4 times pi times radius squared = 4 times 3.14 times 8.6 times 8.6 = 928.9 SA=Surface Area SA=928.9. Volume of a Volley Ball = 4 times pi times r cubed over 3 = 4 times 3.14 times 8.6 over 3 = 36 V=Volume V = 36 cubic inches.

OpenStudy (farmdawgnation):

This doesn't really belong in OpenStudy Feedback. ;)